Considering natural gas in a laboratory burner to be pure methane, CH4, calculate the number of grams of carbon dioxide that would result from the complete burning of 35.0L of methane, measured at 749 torr and 22 degrees.
1
Expert's answer
2018-01-29T13:58:07-0500
CH4 + 2O2 → CO2 + 2H2O

V (CH4) = 35 L

By the law of volumetric relations of Gay-Lussac: V (CH4) = V (CO2)

V (CO2) = 35 L

n (CO2) = V/Vm = 35 L / 22.4 L/mol = 1.5625 mol

M (CO2) = 12 + 16·2 = 44 g/mol

m (CO2) = n·M = 1.5625 mol·44 g/mol = 68.75 g